### v2

# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

Abstract
Video super-resolution aims to enhance low-resolution videos by leveraging both spatial and temporal information. While deep learning has led to impressive progress, it typically requires centralized data, which raises privacy concerns. Federated learning offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs. To address this, we introduce FedVSR, the first FL framework specifically designed for VSR. It is model-agnostic and stateless, and introduces a lightweight loss function based on the DWT to better preserve high-frequency details during local training. Additionally, a loss-aware aggregation strategy combines both DWT-based and task-specific losses to guide global updates effectively. Extensive experiments across multiple VSR models and datasets demonstrate that FedVSR consistently outperforms existing FL methods, achieving up to 0.82 dB higher PSNR, 0.0327 higher SSIM, and 0.0251 lower LPIPS. These results underscore FedVSR’s ability to bridge the gap between privacy and performance, setting a new benchmark for federated learning in low-level vision tasks. The code is available at: https://github.com/alimd94/FedVSR
1 Introduction
Refer to caption (a) Frame 029, Clip 011, REDS[40]
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 30.41, 0.8600, 0.1777 Refer to caption (d) FedProx[27] 30.07, 0.8531, 0.1851
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 26.78, 0.7148, 0.2516 Refer to caption (g) FedVSR (Ours) 32.00, 0.8889, 0.1508
Figure 1: Comparison of RVRT[29] output trained with different federated learning algorithms. General FL methods struggle to reconstruct fine details and textures. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Video super-resolution (VSR) is an advanced computational process that reconstructs high-resolution (HR) video sequences from low-resolution (LR) counterparts, leveraging both spatial and temporal information [32]. Unlike single-image super-resolution (SISR), which relies solely on intra-frame details, VSR exploits temporal correlations across consecutive frames to enhance the quality and sharpness of the output. This capability is crucial in real-world applications such as surveillance [12], video streaming [23], and medical imaging [45, 42], where visual fidelity and temporal coherence are mission-critical. In practice, VSR must also cope with degradations such as down-sampling [60], noise [63], blur [38], and compression artifacts [11]. However, modern deep-learning pipelines address these degradations with data-hungry networks trained on centrally collected data, but centralization itself raises privacy, regulatory, and bandwidth concerns, especially when the raw videos contain sensitive or personally identifiable information.
Federated learning (FL) provides a principled remedy to these concerns [36]. In many VSR scenarios, such as smartphones, surveillance, health-related domains, and remote edge sensors, videos are captured on devices with limited bandwidth and strict privacy constraints. Uploading hours of footage can consume tens of gigabytes of storage; for instance, even a short 10-second 8K video at 30 fps in raw format can exceed 10 GB in size [2]. Additionally, uploading such data may violate policies and laws such as GDPR or HIPAA [19]. Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one. This approach minimizes communication overhead and accommodates the limited computational resources of edge devices. These combined benefits make FL a logical and effective foundation for VSR systems that face such concerns.
Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [21, 27, 51], facial recognition [41, 34], and segmentation [37, 34]. However, their application in super-resolution (SR) remains largely unexplored. Existing studies [39, 56] found that directly applying existing general FL techniques to SR tasks results in poor-quality reconstructions with blurred textures, mainly due to their inability to effectively capture rich low-level information such as texture, edges and motion. Motivated by these limitations, we propose FedVSR, a novel model-agnostic framework for VSR under FL. FedVSR treats VSR models as black boxes, allowing it to work with a variety of architectures without requiring architectural changes or access to intermediate layers. To improve performance under FL constraints, we introduce a lightweight loss term based on the Discrete Wavelet Transform (DWT), which emphasizes high-frequency details that are often lost during decentralized training. While high-frequency preservation is valuable for all SR tasks, its importance is amplified in VSR due to the added complexity of maintaining temporal coherence across frames. In federated settings, data heterogeneity and limited communication intensify this challenge. FedVSR addresses these issues without maintaining state between rounds, making it both scalable and practical for privacy-sensitive, resource-constrained environments. In summary, the main contributions of this paper are as follows:
  • We design a frequency-aware loss using DWT to sharpen fine details during client updates, and introduce a complementary aggregation method to weight model contributions more effectively. To the best of our knowledge, this is the first algorithm specifically designed to address VSR in the FL setting, effectively bridging the gap between privacy preservation and reconstruction quality.
  • Our approach makes no architecture-specific assumptions and does not require storing past weights, making it stateless and model-agnostic. This design reduces memory usage, avoids privacy risks, and eliminates synchronization overhead in federated learning on edge devices.
  • Extensive experiments validate FedVSR’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
2 Related Works
In this section, we begin by discussing the most significant contributions in general FL and VSR. We then explore studies that specifically focus on applying FL to the SR task.
2.1 Federated Learning
Federated learning (FL) is a decentralized learning paradigm that enables multiple clients to collaboratively train models while preserving data privacy. It has been widely applied across diverse domains, including healthcare, finance, and IoT, where data security and regulatory constraints are critical [3, 7, 62]. A foundational approach in this domain is FedAvg [36], which follows a four-step iterative process: (1) the server distributes a global model to all participating clients, (2) each client trains its local model using its own data, (3) the locally updated models are sent back to the central server, and (4) the server aggregates these updates to refine the global model for the next training round.
To enhance the performance of federated learning, multiple methods focus on optimizing local updates. FedProx [27] incorporates a proximal term in the local objective to ensure consistency between local and global models. SCAFFOLD [21] introduces control variates for both clients and the server. FedNova [51] employs normalized averaging to address objective inconsistencies. More recently, MOON [25] applies contrastive learning to assess similarity among different model representations.
2.2 Video Super Resolution
Early VSR methods that applied SISR techniques frame-by-frame neglected the temporal context, resulting in temporal inconsistencies, loss of detail, and undesired smoothing in the high-resolution output [33]. VSR leverages multiple temporally correlated low-resolution frames within a video sequence to generate high-resolution output. Unlike SISR, VSR must account for both spatial and temporal dependencies across frames, making it a highly complex, non-linear, multi-dimensional problem.
Recent advancements in VSR models have introduced various techniques to enhance performance. VRT [30] enables parallel processing of long video sequences using a mutual attention mechanism for feature extraction, alignment, and fusion. RVRT [29] combines parallel and recurrent approaches, employing guided deformable attention for precise clip alignment. IART [54] uses implicit resampling with local cross-attention to preserve frequency details while reducing spatial distortions. DiQP [11], the first model for 8K videos, employs Denoising Diffusion to counteract compression artifacts while maintaining robustness across varying levels.
2.3 Federated Learning on Super Resolution Task
For low-level vision tasks, FedMRI [15] introduces a federated learning framework for MRI reconstruction with a shared encoder for generalized features and client-specific decoders for personalization. FedPR [16] improves image quality using pre-trained models and prompt learning with minimal trainable parameters. p²FedSR [55] enhances privacy-preserving SR by integrating client-specific degradation priors and a selective model aggregation strategy for personalized FL. FedSR [56] employs a detail-assisted contrastive loss for feature alignment at the client level and a hierarchical aggregation policy at the server level with historical global model, optimizing layer-wise model integration.
3 Preliminaries
3.1 VSR Problem Definition
VSR enhances video resolution by using multiple low-resolution frames, leveraging temporal information to exploit inter-frame dependencies, unlike SISR.
Problem Formulation
In addition to the widely used RGB color space, other representations such as the YUV and YCbCr color spaces are frequently employed in VSR applications. Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in a given LR video sequence, while $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ represents its corresponding high-resolution (HR) counterpart, where $s$ is the scaling factor (e.g., $s=2,4,8$). Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius. The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as:
$I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha})$ (1)
where $\phi(\cdot;\cdot)$ is the degradation function parameterized by $\theta_{\alpha}$, incorporating various degradation factors such as noise, motion blur, and down-sampling effects.
A more detailed expression for the degradation process can be written as:
$I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j}$ (2)
where $D$ represents the down-sampling operation, $B$ denotes the blurring effect, $E_{i\to j}$ is a warping function that accounts for motion between frames $\hat{I}_{i}$ and $\hat{I}_{j}$, and $n_{j}$ represents the image noise. The goal of a VSR algorithm is to reconstruct an HR video sequence from degraded LR frames in a manner that closely resembles the ground truth video. This makes this task an ill-posed problem because there exist infinitely many possible HR images $\hat{I}_{i}$ that can produce the same LR image $I_{i}$ after degrading, meaning the degradation operator is highly non-injective [49]. While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence. The inverse process of degradation, which represents the super-resolution task, can be expressed as:
$\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta})$ (3)
where $\tilde{I}_{i}$ is the estimated HR frame corresponding to the ground truth $\hat{I}_{i}$, and $\theta_{\beta}$ represents the model parameters involved in the reconstruction process [33].
3.2 FL Problem Definition
FL enables multiple clients to train a shared model collaboratively without sharing raw data, enhancing privacy and reducing communication overhead through local computations and periodic model aggregation.
Problem Formulation
FL aims to minimize the following objective function across participating clients:
$\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w})$ (4)
where:
  • $K$ is the total number of clients,
  • $p_{k}$ is the relative weight of client $k$ (typically based on its dataset size),
  • $F_{k}(\mathbf{w})$ is the local loss function for client $k$,
  • $\mathbf{w}$ represents the global model parameters.
Each client updates the global model locally:
$\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t})$ (5)
where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme:
$\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}$ (6)
Not all clients participate equally in every training round due to data heterogeneity, resource constraints, network instability, or other limitations [59]. Client averaging based on their contributions stabilizes training, reduces variance, and improves generalization [6].
However, weighting clients solely by dataset size, as in traditional FedAvg[36], does not fully capture their impact on model performance [44]. A client’s importance extends beyond data quantity and is also influenced by the quality of its model updates and the loss it generates during training. Based on [46], we implement an approach where clients with lower loss are typically better trained and can be assigned higher weights during averaging. This ensures improved overall performance by leveraging the robustness properties analyzed in the paper. It is worth noting that some works instead prioritize clients with higher loss to increase fairness [26]. This approach, which contrasts with a focus on final performance, is more common in the context of personalized federated learning.
By incorporating loss-based importance into the aggregation process, federated learning can achieve more reliable and efficient training. This adaptive weighting ensures that client updates more accurately reflect their significance in improving the global model, leading to a more stable and equitable learning process [17].
4 Methodology
VSR remains a challenging problem due to its ill-posed nature. Severe degradations amplify artifacts, making it difficult to balance detail enhancement and noise suppression. To improve generalizability, diverse training data and large batch sizes are needed, but this significantly increases computational costs [8]. FL adds further complexity to VSR due to real-world constraints [14] and the nature of the task. In heterogeneous environments, differences in client data distributions lead to biased updates, poor generalization [5], and inefficient model merging. Gradient mismatches [53] and parameter inconsistencies slow convergence [10]. Most FL algorithms are designed for image classification which can be considered as a well-posed problem [18]. However, VSR is an ill-posed task, requiring a further fundamental assessment of existing methods to ensure effective adaptation. For instance, FedProx assumes a certain degree of similarity across client data distributions, but in real-world applications, these distributions can vary significantly. When this assumption fails, models experience drift, leading to inconsistent local updates and poor generalization [13]. Similarly, SCAFFOLD assumes unbiased client updates, yet in large-scale deployments, heterogeneous data causes significant divergence, slowing convergence. Moreover, its need for additional state information per client makes it impractical for large-scale FL systems due to communication and storage constraints [10]. Meanwhile, Other methods such as FedDisco[58] and MOON rely on class label distributions or logits, making them inherently task-specific rather than generalizable.This work proposes a solution to enhance the overall objective of VSR models while improving aggregation to mitigate these challenges.
4.1 Enhancing Loss Function for Federated VSR
Selecting a proper loss function as objective for VSR model is crucial and determine the overall output’s performance [50, 47] and as depicted in Fig.1. It is evident that existing FL algorithms has problems in capturing and generating details in the output. Since we aim to avoid adding computational overhead to the client by using adversarial or perceptual loss, and because our objective is to keep the model untouched, ensuring FedVSR remains model-agnostic, we introduce an additional term to the overall loss of model improve performance.
4.1.1 DWT for High-Frequency Preservation
One common failure in training VSR models is producing overly smooth outputs, meaning the trained model fails to capture high-frequency information. As a result, details and textures are missing from the output. Since the models we used are among the top-performing ones in a centralized setting, this suggests that the failure to restore high-frequency details is specifically related to training under the FL paradigm. To alleviate this, we leverage the Discrete Wavelet Transform (DWT), which has proven effective in super-resolution tasks [20, 61].
DWT is a multi-resolution signal decomposition technique that represents a signal in terms of approximation and detail coefficients at different scales. For a 2D signal (such as an image or video frame), DWT decomposes the signal into sub-bands of different frequencies and orientations such as:
  • Approximation Coefficients $F_{i}^{LL}$ (low-pass filtered in both dimensions).
  • Horizontal Detail Coefficients $F_{i}^{LH}$ (high-pass in vertical, low-pass in horizontal).
  • Vertical Detail Coefficients $F_{i}^{HL}$ (low-pass in vertical, high-pass in horizontal).
  • Diagonal Detail Coefficients $F_{i}^{HH}$ (high-pass in both dimensions).
One of the simplest wavelets used in DWT is Haar wavelet, which consists of a step function. It is computationally efficient and commonly used in image and video processing tasks. In this method each frame’s pixel values are convolved with:
  • Low-pass filter $h=\frac{1}{\sqrt{2}}[1,1]$ to extract approximation coefficients.
  • High-pass filter $g=\frac{1}{\sqrt{2}}[1,-1]$ to extract detail coefficients.
This process is repeated independently for each frame in the sequence.
The coefficient in those four sub-bands are calculated as:
$F_{i}^{LL}=(F_{i}*h^{T})*h,\quad F_{i}^{LH}=(F_{i}*h^{T})*g$ (7)
$F_{i}^{HL}=(F_{i}*g^{T})*h,\quad F_{i}^{HH}=(F_{i}*g^{T})*g$ (8)
The low-frequency component $F_{i}^{LL}$, contains the content and color information of the input frame, while the three high-frequency sub-bands capture detailed information about global structures and textures [43]. Therefore, for our loss helper term, we consider only the high-frequency sub-bands and ignore the low-frequency ones.
The loss is then computed pixel-wise in each sub-band between the predicted $\bar{F}$ and ground truth $\hat{F}$ frame at time step $i$:
$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH}$ (9)
$L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL}$ (10)
$L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (11)
Then with the Charbonnier loss [9] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ :
$\mathcal{L}_{\text{HiFr}}=\sum_{i}\sum_{s\in\{LH,HL,HH\}}\sqrt{{L_{i}^{s}}^{2}+\epsilon^{2}}$ (12)
Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows:
$\mathcal{L}_{\text{total}}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$ (13)
As our experiments demonstrate (see Section 5.3), the optimal weight values are unbounded, meaning they are not constrained by the relationship $\lambda_{\text{VSR}}=1-\lambda_{\text{HiFr}}$ and can be independently adjusted for better performance.
4.2 Loss-Aware Aggregation for Federated VSR
To further enhance the model’s performance, we employ a greedy aggregation approach with a decay rate, inspired by MKL-SGD [46]. MKL-SGD improves robustness by selecting samples with the lowest loss during training, thereby mitigating the impact of outliers. Similarly, in our method, clients accumulate their loss values throughout local training and compute an average loss based on the number of training steps. When sending their updated model weights to the server, they also transmit this average loss. At the server, we perform inverse weighted averaging, where each client’s weight is computed as $I_{i}=\frac{1}{\mathcal{L}_{i}}$ and adjusted using an adaptive step: $I^{\prime}_{i}=I_{i}^{\alpha}$ where $\alpha$ controls the strength of loss-based weighting. To regulate this process, we introduce a decay rate that gradually reduces the effect of loss-based weighting over time, following:$\alpha=\alpha^{1-\frac{t}{T}}$ where $t$ is the current round and $T$ is the total number of rounds. This ensures that early training benefits from loss-based prioritization, while later rounds transition smoothly toward standard averaging, balancing robustness and generalization. The overall algorithm is summarized in Algorithm 1.
Algorithm 1 Our Method
1:Number of communication rounds $T$, Number of clients $N$, Learning rate $\eta$, Adaptive step $\alpha$, Decay rate $\gamma$
2:Global model weights $W_{T}$
3:Server Initialization Initialize global model $W_{0}$
4:for $t=1,\dots,T$ do
5:    Server selects a subset of clients randomly $S_{t}\subseteq\{1,\dots,N\}$
6:    for each client $i\in S_{t}$ in parallel do
7:        Client updates local model:
8:         Retrieve latest global model $W_{t}$
9:         Perform local training for $E$ epochs with learning rate $\eta$
10:        Compute Local Loss:
11:         Compute model-specific loss $\mathcal{L}_{\text{VSR}}$
12:         Compute $\mathcal{L}_{\text{DWT}}$ loss $\mathcal{L}_{\text{DWT}}$
13:         Compute total loss
$\mathcal{L}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$
14:        Client updates local model:
$\mathbf{w}_{i}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{i}(\mathbf{w}^{t})$
15:         Return updated weights $W_{i}$ and loss value $\mathcal{L}_{i}$ to server
16:    end for
17:    Server Aggregation: (Performed at the Server)
18:     Compute inverse loss weighting $I_{i}=\frac{1}{\mathcal{L}_{i}}$
19:     Apply adaptive step $I^{\prime}_{i}=I_{i}^{\alpha}$
20:     Normalize inverse loss weights $w_{i}=\frac{I^{\prime}_{i}}{\sum_{j\in S_{t}}I^{\prime}_{j}}$
21:     Compute weighted model update $W_{i}^{\text{scaled}}=w_{i}\cdot W_{i}$
22:     Aggregate global model $W_{t+1}=\sum_{i\in S_{t}}W_{i}^{\text{scaled}}$
23:    Apply Decay Rate to Adaptive Step $\alpha=\alpha^{1-\frac{t}{T}}$
24:end for
25:return Global model $W_{T}$
5 Experiments
Model Algorithm Less Heterogeneous More Heterogeneous
REDS[40] Vid4[31] REDS[40] Vid4[31]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
VRT [30] FedAvg[36] 29.71 0.8522 0.1916 24.88 0.7680 0.2446 28.85 0.8282 0.2200 24.60 0.7516 0.2557
FedProx[27] 28.82 0.8281 0.2133 24.18 0.7341 0.2668 28.64 0.8236 0.2222 24.10 0.7256 0.2713
SCAFFOLD[21] 28.59 0.8181 0.2228 23.99 0.7182 0.2737 28.17 0.8074 0.2318 24.04 0.7208 0.2739
FedVSR (Ours) 29.86 0.8562 0.1908 24.92 0.7682 0.2393 29.03 0.8339 0.2123 24.73 0.7584 0.2492
RVRT [29] FedAvg[36] 31.30 0.8877 0.1536 25.96 0.8156 0.1974 29.89 0.8578 0.1867 25.67 0.8040 0.2065
FedProx[27] 30.09 0.8618 0.1899 25.20 0.7775 0.2330 29.59 0.8526 0.1943 25.02 0.7730 0.2353
SCAFFOLD[21] 29.63 0.7967 0.2101 24.79 0.7317 0.2407 26.71 0.7142 0.2590 23.08 0.6441 0.3018
FedVSR (Ours) 31.54 0.8922 0.1455 26.05 0.8191 0.1889 30.90 0.8811 0.1576 25.90 0.8131 0.1910
IART [54] FedAvg[36] 30.56 0.8729 0.1695 25.43 0.7949 0.2118 29.52 0.8493 0.1909 25.04 0.7774 0.2272
FedProx[27] 29.65 0.8500 0.1683 24.72 0.7599 0.2307 29.30 0.8444 0.1958 24.42 0.7505 0.2293
SCAFFOLD[21] 30.29 0.8678 0.1692 25.18 0.7857 0.2084 29.26 0.8413 0.1992 24.85 0.7690 0.2278
FedVSR (Ours) 30.88 0.8812 0.1556 25.54 0.8012 0.1990 29.76 0.8536 0.1887 25.26 0.7850 0.2220
Table 1: PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$) comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
5.1 setup
Dataset:
Following [30, 29, 54], one of the most commonly used datasets for training VSR is REDS [40]. It contains 240, 30, and 30 sequences for training, validation, and testing, respectively, with each sequence consisting of 100 frames at a resolution of 720 × 1280. The corresponding low-resolution (LR) frames are generated through bicubic downsampling from the high-resolution (HR) images. This dataset features a variety of natural and handmade objects, locations, and diverse content, making it heterogeneous in decentralized settings [40]. To further evaluate FedVSR and assess its robustness in more realistic, highly heterogeneous scenarios, we conducted additional experiments using Kinetics-400 [22], a dataset primarily used for human action recognition. It includes 400 action classes with 306,245 videos, each approximately 10 seconds long, sourced from unique YouTube videos. Since Kinetics-400 contains videos of varying resolutions, we selected only those with a 720 × 1280 resolution to ensure a fair comparison with REDS. From the 400 classes, we randomly chose 40, and for each class, selected six videos, extracting 100-frame clips to match REDS’ format. Kinetics-400 is particularly well-suited for testing FedVSR under more extreme heterogeneity because, unlike professionally filmed datasets, its videos often exhibit camera shake, lighting variations, shadows, and background clutter [22]. This makes it a challenging but realistic benchmark for assessing the model’s adaptability in diverse and less controlled environments. For consistency with our experimental setup, we refer to REDS as the ”less heterogeneous” dataset and Kinetics-400 as the ”more heterogeneous” dataset, given the increased variability in content, motion, and environmental conditions in the latter.
Implementation Details:
To maintain our objective of developing a model-agnostic framework, we did not alter the official parameters of the baseline VSR models. Instead, we used their recommended settings for training. We created 40 clients, each assigned six videos in both experimental setups. In the more heterogeneous setup, each client received videos from a single class; for example, one client had only ”archery” videos, while another had only ”making pizza” videos. In contrast, in the less heterogeneous setup, the six videos were randomly assigned to each client. We set the total number of global rounds to 100, with each client performing one local round (epoch) per global round. In each round, we randomly selected four clients to perform training. After completing their local training, these clients sent their updated weights to the server for aggregation. All experiments were conducted using bi-cubic (BI) degradation.
Baselines:
For our baseline, we use a combination of FedAvg[36], FedProx[27], and SCAFFOLD[21] from the FL side, along with VRT[30], RVRT[29], and IART[54] from VSR. We selected these FL methods as they are widely accepted and frequently used to benchmark core algorithmic performance in recent literature. While newer approaches like MOON[24], FedDyn[1], FedDisco[57], FedLAW[28] exist, they are rely on task-specific assumptions such as logits, class-wise statistics, or server-side data proxies that do not translate well to low-level, regression-based tasks like VSR. On the other hand, exisitng FL-SR models such as FedSR, p²FedSR, and FedMRI are not directly comparable to FedVSR for two key reasons. First, they are tightly coupled to a specific model architecture: they require dividing the model into predefined blocks (“shallow–deep–upsampler”), which limits flexibility and demands structural changes to existing SR models. Second, these methods are stateful meaning each client must store additional data across training rounds, such as degradation descriptors, deep feature extractors, or previous model weights. This leads to significantly higher storage, computation, and coordination overhead. In contrast, FedVSR is designed to be both model-agnostic and stateless, making it simpler, more lightweight, and easier to apply in diverse real-world scenarios.
Evaluation Metrics:
The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS).
Figure 2: PSNR values across different rounds for various test sets under different settings for VRT [30], RVRT [29], and IART [54].
Reproducibility:
To ensure our experiments are reproducible, we fixed the random seed, which doesn’t affect client selection, model initialization, etc. We used four NVIDIA A6000 GPUs for training, with each model under a federated learning algorithm taking an average 46 hours to train. For the baselines alone, this consumed 36 days (2 experiments × 9 models+algorithms × 2 days per run).
5.2 Results and Comparison
Quantitative Results:
As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets. On average, FedVSR surpasses existing FL methods by a margin of 0.61 dB in PSNR and 0.022 in SSIM and 0.022 in LPIPS in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.66 dB in PSNR and 0.028 in SSIM and 0.019 in LPIPS, demonstrating its robustness under challenging conditions. Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin. In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods.
The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results. As shown in Fig.2, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FedAvg, FedProx, and SCAFFOLD, particularly in later training rounds. This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions. The study suggests that FedProx’s limiting points may remain far from the true minimizer of the global empirical risk [48], which could explain its suboptimal performance in our experiments. Although FedProx may still generalize well by aggregating local data effectively in certain conditions, its weaker convergence properties become apparent in more heterogeneous scenarios. We also observe that FedProx is highly sensitive to the $\mu$ value during training. The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases. However, we notice that for any $\mu$ greater than 0.001, the VSR model struggles significantly with learning. This suggests that the relationship is more complex and requires further investigation and analysis.
Similarly, our findings regarding SCAFFOLD’s inefficacy in highly heterogeneous settings corroborate previous research that highlights its susceptibility to instability under heterogeneous data [4]. The significant performance drops and stagnation observed in our experiments, particularly with the RVRT model, align with the reported issues of SCAFFOLD transmitting nearly twice the data of other algorithms due to its reliance on control variates. This added communication burden does not necessarily translate to better convergence, as our results show that SCAFFOLD struggles to optimize effectively in diverse settings. Prior work [4] also notes that SCAFFOLD is highly prone to sudden performance breakdowns when gradient clipping is not employed, which is reflected in our results where its PSNR fluctuates or plateaus early in highly heterogeneous scenarios. The increased variance in SCAFFOLD’s performance under skewed data distributions further supports our findings that it fails to adapt smoothly across different client datasets, reinforcing its limitations in federated VSR tasks.
Overall, these existing studies provide strong theoretical support for our experimental results. In contrast, FedVSR’s convergence, stability, and robustness across varying levels of data heterogeneity highlight its effectiveness in overcoming these limitations.
Qualitative Results:
As showen in Fig. 1 clearly demonstrate the limitations of FedAvg, FedProx, and SCAFFOLD in accurately reconstructing fine details. While these methods reduce some pixelation from the LQ input, they all fail to recover the textures and sharp edges present in the ground-truth. FedAvg produces overly smooth results, blurring important structural details such as the wooden lattice and frame edges. FedProx exhibits similar blurring, struggling to restore high-frequency textures. SCAFFOLD, not only fails to produce a sharp reconstruction but also introduces noticeable artifacts, distorting the texture of the wooden lattice and making the output appear unnatural. In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches. Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 8.
5.3 Ablation Study
Evaluating $\mathcal{L}_{\text{HiFr}}$ Loss and Adaptive Aggregation
To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments. Using FedAvg as a baseline, we added each component separately to evaluate their individual effects. To assess the impact of the $\mathcal{L}_{\text{HiFr}}$ term, we applied it to FedAvg and observed the changes (F-$\mathcal{L}_{\text{HiFr}}$). Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds. We then compared this approach to an adaptive version that incorporates a decay mechanism (F-ALA). As shown in Table 2, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases. Additionally, both methods individually surpass FedAvg [36]. However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics. These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights. Due to space limitations, LPIPS metric results are provided in the supplementary material 7
Model Setting Less Heterogeneous More Heterogeneous
REDS Vid4 REDS Vid4
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 24.91 0.7681 28.98 0.8322 24.70 0.7577
F-ALA 29.75 0.8531 24.89 0.7681 28.90 0.8313 24.64 0.7559
F-GLA 29.70 0.8523 24.86 0.7670 28.83 0.8279 24.55 0.7500
RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 26.02 0.8178 30.24 0.8617 25.84 0.8042
F-ALA 31.38 0.8880 26.00 0.8171 30.19 0.8601 25.79 0.8041
F-GLA 31.31 0.8879 25.95 0.8155 29.88 0.8576 25.60 0.8032
IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 25.46 0.8001 29.64 0.8508 25.21 0.7837
F-ALA 30.70 0.8789 25.44 0.7998 29.60 0.8500 25.18 0.7831
F-GLA 30.57 0.8730 25.42 0.7951 29.50 0.8482 24.99 0.7761
Table 2: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR and SSIM under different heterogeneity settings.
Optimal $\lambda$s
Using a grid search, we explored different values for $\lambda_{\text{VSR}}$ and $\lambda_{\text{HiFr}}$ to determine the optimal ones. We varied each parameter from 1 to 0.1 in steps of 0.1 under both bounded and unbounded conditions. Due to the high computational cost of the experiments, we trained the VSR models for only 20 rounds and selected the best-performing configurations. Our results show that for VRT[30], the optimal values for both parameters are 1 For RVRT[29], the best values are 1 and 0.5, while for IART[54], the optimal values are 0.5 and 0.5, respectively.
Empirical Fisher for Estimating Weights Importance
Inspired by [35], we explored estimating the importance of each weight in client models to enhance aggregation using Empirical Fisher (EF) Information. EF approximates the Fisher information matrix by leveraging gradients collected during backpropagation after training. While this method led to slight performance improvements, it proved computationally too expensive for clients, making it impractical for our setting. A more efficient variant of EF was introduced in [52]; however, as it is tailored for classification tasks, we leave its adaptation to VSR for future work.
6 Conclusion
In this work, we introduced FedVSR, the first FL framework specifically designed for VSR, addressing privacy concerns while preserving high-quality performance. Unlike existing FL approaches that struggle with low-level vision tasks, our model-agnostic and stateless design ensures adaptability across various VSR architectures without prior model knowledge or historical weight storage. By incorporating a lightweight loss term, we effectively guide local updates and enhance global aggregation, resulting in notable performance improvements over general FL baselines. Our method functions as a plug-and-play solution, and extensive experiments confirm its superiority over existing FL approaches, demonstrating its potential for real-world, privacy-sensitive applications. Furthermore, our study highlights key challenges in applying FL to low-level vision tasks, offering valuable insights for future research in this under-explored domain. As the first FL-based approach for VSR, this work establishes a strong baseline and provides a clear research path for future advancements. Despite promising results, a gap remains between FL-based VSR and centralized approaches. Future research could refine the aggregation strategy with a more intelligent, task-specific approach tailored for low-level vision. Additionally, our loss-guided optimization could be further enhanced to improve stability and better address existing challenges in FL-based VSR.
Supplementary Material
7 Extended Evaluation of HiFr Loss and Adaptive Aggregation Using LPIPS
To complement the analysis based on PSNR and SSIM discussed earlier in the paper, we present here an additional evaluation using LPIPS, which serves as a perceptual similarity metric. This acts as an extension to the previously discussed metrics and provides a more complete picture of model performance. Due to space limitations, we include this analysis here.
To evaluate the impact of the DWT loss term and the loss-aware adaptive aggregation strategies, we conducted a set of ablation experiments using FedAvg as the baseline. Each component was added independently to isolate its individual effect. Specifically, we assessed the contribution of the $\mathcal{L}_{\text{HiFr}}$ term by applying it to FedAvg (F-$\mathcal{L}_{\text{HiFr}}$), and examined two aggregation variants: a greedy loss-aware approach (F-GLA) that prioritizes updates from clients with lower losses, and an adaptive version (F-ALA) that incorporates a decay factor to regulate client influence over time.
As shown in Table 3 with LPIPS results, the observed trends remain largely consistent across metrics. The DWT loss term consistently provides more significant gains compared to the aggregation strategies, and both methods individually outperform the standard FedAvg baseline. However, similar to PSNR and SSIM, the greedy strategy sometimes leads to only marginal improvements or even minor performance degradation, especially under high heterogeneity. These findings underscore the importance of high-frequency information not only for boosting perceptual quality but also for reliably identifying valuable client updates.
Model Setting Less Heterogeneous More Heterogeneous
REDS Vid4 REDS Vid4
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 0.1910 24.91 0.7681 0.2399 28.98 0.8322 0.2130 24.70 0.7577 0.2500
F-ALA 29.75 0.8531 0.1912 24.89 0.7681 0.2411 28.90 0.8313 0.2140 24.64 0.7559 0.2520
F-GLA 29.70 0.8523 0.1914 24.86 0.7670 0.2436 28.83 0.8279 0.2198 24.55 0.7500 0.2555
RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 0.1506 26.02 0.8178 0.1910 30.24 0.8617 0.1612 25.84 0.8042 0.1978
F-ALA 31.38 0.8880 0.1518 26.00 0.8171 0.1931 30.19 0.8601 0.1728 25.79 0.8041 0.1992
F-GLA 31.31 0.8879 0.1527 25.95 0.8155 0.1948 29.88 0.8576 0.1807 25.60 0.8032 0.2011
IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 0.1588 25.46 0.8001 0.1999 29.64 0.8508 0.1899 25.21 0.7837 0.2231
F-ALA 30.70 0.8789 0.1591 25.44 0.7998 0.2101 29.60 0.8500 0.1903 25.18 0.7831 0.2246
F-GLA 30.57 0.8730 0.1690 25.42 0.7951 0.2119 29.50 0.8482 0.1929 24.99 0.7761 0.2284
Table 3: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR, SSIM, and predicted LPIPS under different heterogeneity settings.
8 Further Analysis: Performance Metrics and Visual Comparisons
This section provides additional performance evaluations for our proposed method. Figures 3, 5, and 7 illustrate the PSNR, SSIM and LPIPS trends across different rounds for multiple test sets under various configurations of VRT [30], RVRT [29], and IART [54]. These results further highlight the effectiveness of our approach in enhancing the performance across different models.
Figures 4, 6, and 8 provide qualitative visual comparisons on various test frames. Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR). The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
Figure 3: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for VRT.
Refer to caption (a) Frame 000, Clip 015, REDS
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 32.41, 0.8981, 0.1456 Refer to caption (d) FedProx[27] 32.10, 0.8935, 0.1511
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 28.85, 0.7625, 0.1934 Refer to caption (g) FedVSR (Ours) 33.79, 0.9169, 0.1151
Figure 4: Comparison of RVRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Figure 5: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for RVRT.
Refer to caption (a) Frame 040, Clip 1, Vid4
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 21.15, 0.7185, 0.3520 Refer to caption (d) FedProx[27] 20.43, 0.6735, 0.4101
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 20.41, 0.6679, 0.3803 Refer to caption (g) FedVSR (Ours) 21.19, 0.7200, 0.3433
Figure 6: Comparison of VRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Figure 7: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for IART.
Refer to caption (a) Frame 070, Clip 000, REDS
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 27.87, 0.8211, 0.2228 Refer to caption (d) FedProx[27] 27.54, 0.8132, 0.2247
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 27.69, 0.8151, 0.2207 Refer to caption (g) FedVSR (Ours) 28.07, 0.8319, 0.2014
Figure 8: Comparison of IART output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.